OpenStudy (anonymous):

if 1/2AB = 1/2EB then AB = EF. what property is it? @vzfreakz

OpenStudy (anonymous):

Multiplicative Property? Not putting a positive twist into this.

OpenStudy (anonymous):

@vzfreakz the only thing that is similar to (multiplicatve property) is (multiplication and division properties) is it that one?

OpenStudy (anonymous):

@surjithayer that is not in the awnser choices

OpenStudy (anonymous):

i am thinking of geometry can you give the choices.

OpenStudy (anonymous):

If it's multiplicative, it probably be multiplication properties. But let's not go straight to the answer. Let's break it down a bit.

OpenStudy (anonymous):

@surjithayer yes it is geometry and its (1/2AB=1/2EF then AB=EF i typed the 'b" on accident idk if that helps a bit

OpenStudy (anonymous):

My official answer that I'm going with would be multiplicative inverse.

OpenStudy (anonymous):

@vzfreakz @surjithayer transitive property reflexive property symmetric property addition and subtraction properties multiplication and divison properties distributive property substitution property

OpenStudy (anonymous):

@vzfreakz those are the awnser choices

OpenStudy (anonymous):

I would go with multiplication based on what I said earlier, multiplicative inverse.

OpenStudy (anonymous):

it is multiplicative and division property if we multiply each side by 2 we get the solution.

OpenStudy (anonymous):

@vzfreakz @surjithayer if 1/2AB=1/2CD which property can be used to justify the statement AB = CD

OpenStudy (anonymous):

same

OpenStudy (anonymous):

Multiplication and Division Property.

OpenStudy (anonymous):

Well, probably not division but you could verify it with @surjithayer.

OpenStudy (anonymous):

Which postulate states that if B is between A and C, then AB + BC = AC? a. ruler postulate c. whole segment postulate b. addition property of congruence d. segment addition postulate @vzfreakz

OpenStudy (anonymous):

The answer could be D.

OpenStudy (anonymous):

Make sure to post your new questions in a new question section.

OpenStudy (anonymous):

@vzfreakz

OpenStudy (anonymous):

@vzfreakz

OpenStudy (anonymous):

and it says " For the proof shown, provide statement 3." @surjithayer

OpenStudy (anonymous):

@trruhgnh hey, just to let you know, that you owe me a medal on your last question.

OpenStudy (anonymous):

@vzfreakz okay there it is, can you help in those images i posted here?

OpenStudy (anonymous):

Give me a minute. I'm working on your other assignments. I'm all over the places at the moment. I see the pictures but did you provide any questions to it>

OpenStudy (anonymous):

@vzfreakz yes it just says for the proof shown provide statement 3